OpenStudy (anonymous):

for f(x)=5x^2, find (f(x+h)-f(x))/h

OpenStudy (anonymous):

go slow first off \(f(x+h)=5(x+h)^2=5(x^2+2xh+h^2)=5x^2+10xh+5h^2\)

OpenStudy (anonymous):

then subtract \(f(x)=5x^2\) leaving you just with \[10xh+5h^2\]

OpenStudy (anonymous):

finally divide each term by \(h\) and you are done

OpenStudy (anonymous):

would the answer be 10x+5h?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

....

OpenStudy (anonymous):

yes